OpenStudy (anonymous):

does (x+1)=(x+1)/(x-1) have any false solutions

hero (hero):

Well basically, we can rewrite it as two fractions: \(\frac{x+1}{1}=\frac{x+1}{x-1}\) If the numerator is the same, I would think the denominator would also have to be the same in order for both sides to be equal. The only thing that would work is x = 2. Anything else would make it false.